# Minimal Torsion Curves in Geometric Isogeny Classes

### v1

@@ Acknowledgements
We thank John Cremona, Álvaro Lozano-Robledo, Jeremy Rouse, Parker Schwartz, and Drew Sutherland for helpful conversations. We thank John Cremona and Filip Najman for helpful comments on an earlier version of this paper. All authors were supported by NSF grant DMS-2137659. The first author was partially supported by an A. J. Sterge Faculty Fellowship and NSF grant DMS-2145270.
@@ 2. Background and Notation
Throughout, $F$ denotes a number field and $\overline{F}$ denotes a fixed algebraic closure. We write $\operatorname{Gal}_{F}$ for the absolute Galois group $\operatorname{Gal}(\overline{F}/F)$.
@@
For a prime number $\ell$, our notation for subgroups of $\operatorname{GL}_{2}(\mathbb{Z}/\ell\mathbb{Z})$ and $\operatorname{GL}_{2}(\mathbb{Z}_{\ell})$ follows [39] and [27], respectively. For $\ell$-adic images, this is known as the “RSZB label” and has the form N.i.g.n, where N is the level, i is the index, g is the genus, and n is a positive integer used to distinguish nonconjugate subgroups. We also refer to specific elliptic curves over $\mathbb{Q}$ by their $L$-functions and Modular Forms Database (LMFDB) label.
@@ 2.2. Galois Representations
which encodes the Galois action on all points in $E(\overline{F})$ of order a power of $\ell$. If $E$ is non-CM, then $\operatorname{im}\rho_{E,\ell^{\infty}}$ is an open subgroup of $\operatorname{GL}_{2}(\mathbb{Z}_{\ell})$ by Serre’s Open Image Theorem [34]. Thus there exists a nonnegative integer $d$ such that $\operatorname{im}\rho_{E,\ell^{\infty}}=\pi^{-1}(\operatorname{im}\rho_{E,\ell^{d}})$, where $\pi:\operatorname{GL}_{2}(\mathbb{Z}_{\ell})\rightarrow\operatorname{GL}_{2}(\mathbb{Z}/\ell^{d}\mathbb{Z})$ is the natural reduction map. The smallest such $\ell^{d}$ for which this holds is called the level of the $\ell$-adic Galois representation.
@@
Aside from $\operatorname{GL}_{2}(\mathbb{Z}/\ell\mathbb{Z})$, all groups which are known to occur as the image of the mod $\ell$ Galois representation associated to a non-CM elliptic curve over $\mathbb{Q}$ appear (up to conjugacy) in Tables 3 and 4 of [39]. This list is complete for $\ell\leq 13$ by work of Zywina [41] and Balakrishnan, Dogra, Müller, Tuitman, and Vonk [1], and it has been conjectured to be complete for all $\ell$ by both Sutherland [39, Conjecture 1.1] and Zywina [41, Conjecture 1.12]. Unconditionally, we have the following result for $\ell\geq 17$.
Theorem 2.1 (Mazur [31], Serre [35], Bilu, Parent, and Rebolledo [2]).
@@
Suppose $E/\mathbb{Q}$ is a non-CM elliptic curve and $\ell\geq 17$ is prime. If $\operatorname{im}\rho_{E,\ell}$ is not equal to $\operatorname{GL}_{2}(\mathbb{Z}/\ell\mathbb{Z})$ and not conjugate to a group in Table 3 or 4 of [39], then $\operatorname{im}\rho_{E,\ell}$ is contained in $C_{ns}^{+}(\ell)$, the normalizer of a non-split Cartan subgroup of $\operatorname{GL}_{2}(\mathbb{Z}/\ell\mathbb{Z})$.
@@
Refinements of Theorem 2.1 appear in [41, Proposition 1.13], which is also proven in [29, Appendix B]. Recently, Furio and Lombardo [19] have shown that proper subgroups of the normalizer of a non-split Cartan subgroup do not occur for primes larger than 37. Taken together with prior work, [19, Theorem 1.6] implies the following result.
Theorem 2.2 (Furio, Lombardo [19]).
@@
Suppose $E/\mathbb{Q}$ is a non-CM elliptic curve and $\ell\geq 17$ is prime. If $\operatorname{im}\rho_{E,\ell}$ is not equal to $\operatorname{GL}_{2}(\mathbb{Z}/\ell\mathbb{Z})$ and not conjugate to a group in Table 4 of [39], then $\operatorname{im}\rho_{E,\ell}$ is conjugate to $C_{ns}^{+}(\ell)$.
@@
Many partial classification results exist for the image of the $\ell$-adic Galois representation of an elliptic curve $E/\mathbb{Q}$. First suppose $E$ is non-CM. The groups which occur as $\operatorname{im}\rho_{E,2^{\infty}}$ are known due to work of Rouse and Zureick-Brown [33]. For odd primes $\ell$, Sutherland and Zywina [40] have identified the images that occur infinitely often, and work of Rouse, Sutherland, and Zureick-Brown [27] provides additional classification results for $3\leq\ell\leq 11$ which are complete up to computing rational points on 6 remaining modular curves. If $E$ has complex multiplication, see work of Lozano-Robledo [30].
@@ 2.3. Elliptic Curves with an Isogeny
If the image of $\chi$ lands in {±$\{\pm 1\}$, then there is a twist of $E$ for which the point corresponding to $P$ becomes $F$-rational. In general, we have the following proposition.
@@
Let $N\geq 3$ be an integer, and let $E/F$ be an elliptic curve with an $F$-rational cyclic isogeny of degree $N$. There is an extension $L/F$ with $[L:F]\mid\frac{\varphi(N)}{2}$ and a quadratic twist $E^{\prime}$ of $E/L$ such that $E^{\prime}(L)$ has a point of order $N$.
@@
This is a consequence of [5, Theorem 5.5]. ∎
@@ 2.4. Modular Curves
In this paper, we are interested in characterizing degrees of points on the modular curve $X_{1}(N)$, where $N$ is a positive integer. Recall $X_{1}(N)$ is an algebraic curve over $\mathbb{Q}$ whose non-cuspidal points correspond to isomorphism classes of elliptic curves with a distinguished point of order $N$. If $E$ is an elliptic curve defined over a number field $F$ with $P\in E(F)$ of order $N$, then $(E,P)$ gives an $F$-valued point on $X_{1}(N)$ via this moduli interpretation. By definition, this is a morphism of $\mathbb{Q}$-schemes $f:\text{Spec}\,F\rightarrow X_{1}(N)$, and the image of $f$ is the associated closed point, denoted $[E,P]$. See [17, Section 7.7], [16], [36, §$\S 6.7$], [37, Appendix C, §$\S 13$], or [15] for more details. If $x\in X_{1}(N)$ is closed point, we define the degree of $x$ to be the degree of the residue field $\mathbb{Q}(x)$. For a non-cuspidal point $x$, we can construct $\mathbb{Q}(x)$ explicitly via the following result.
@@
See, for example, [8, Lemma 2.5], and [15, p. 274, Proposition VI.3.2]. ∎
@@
If $E/\mathbb{Q}(j(E))$ corresponds to an equation of the form $y^{2}=x^{3}+Ax+B$ and $P=(x_{0},y_{0})\in E$, then we may take
@@
Note $AB\neq 0$ if $E$ is non-CM. Thus by Lemma 2.4 we can compute the degree of a closed point on $X_{1}(N)$ associated to a non-CM elliptic curve by factoring division polynomials. See [36, p. 107] for details, including a formulation of the Weber function which is more clearly model-independent.
@@
The moduli interpretation ensures the map is defined over $\mathbb{Q}$, and the degree calculation follows from [17, p.66]. ∎
@@ 2.5. Complex Multiplication
where $\Delta_{K}$ is the discriminant of $K$. We have #$\#\mathcal{O}^{\times}=2$ unless $\Delta=-3$ or $-4$ in which case #$\#\mathcal{O}=6$ or $4$, respectively. If $E$ has CM by the order $\mathcal{O}$ in $K$, then $[\mathbb{Q}(j(E)):\mathbb{Q}]=h(\mathcal{O})$ by [13, Theorem 11.1], the class number of $\mathcal{O}$. If $\mathfrak{f}=1$ then $h(\mathcal{O})=h_{K}$, the class number of $K$. If $\mathfrak{f}>1$, then by [13, Corollary 7.24]
(1) $h(\mathcal{O})=[\mathbb{Q}(j(E)):\mathbb{Q}]=h_{K}\frac{2}{w_{K}}\mathfrak{f}\prod_{p\mid\mathfrak{f}}\left(1-\left(\frac{\Delta_{K}}{p}\right)\frac{1}{p}\right).$
@@ 3.1. Fields of Definition for Isogenies
Let $\mathcal{E}$ be a rational $\overline{\mathbb{Q}}$-isogeny class of non-CM elliptic curves. By definition, there exists $E_{0}\in\mathcal{E}$ with $j(E_{0})\in\mathbb{Q}$, and for any $E\in\mathcal{E}$ there is a $\overline{\mathbb{Q}}$-isogeny $\varphi:E\rightarrow E_{0}$. Since the degree closed points on $X_{1}(N)$ can be computed using any Weierstrass model of $E$ by Lemma 2.4, we are free to replace $E$ and $E_{0}$ by quadratic twists in order to achieve a more convenient representation of $\varphi$. The lemma given below essentially follows from the fact that $\mathbb{Q}(j(E),j(E_{0}))=\mathbb{Q}(j(E))$ is contained in the residue field of any closed point on $X_{1}(N)$ associated to $E$, and this is the field of moduli of the isogeny $\varphi$; see [11, §3.3§3.3$\S$ 3.3§ 3.3] or [14, Corollary A.5].
@@
Since $\mathcal{E}$ is rational, there exists $E_{0}\in\mathcal{E}$ with $j(E_{0})\in\mathbb{Q}$. By definition there exists an isogeny $\varphi:E\rightarrow E_{0}$ defined over $\overline{\mathbb{Q}}$ which we may assume is cyclic of degree $N$; see Lemma A.1 in [14]. Let $C$ denote its kernel. Note that $F=\mathbb{Q}(j(E),\mathfrak{h}(P))$ by Lemma 2.4 and there exists a Weierstrass equation of $E/F$ with $P\in E(F)$. The proof of [11, Proposition 3.2] shows $C$ is $F$-rational, as we will now show. Suppose $\sigma(C)\neq C$ for some $\sigma\in\operatorname{Gal}_{F}$, and consider the induced isogeny $E^{\sigma}\rightarrow(E/C)^{\sigma}$. Since $j(E/C)=j(E_{0})\in\mathbb{Q}$, we see that $j((E/C)^{\sigma})=j(E_{0})$. Thus composition with an isomorphism to $E_{0}$ yields a cyclic $N$-isogeny $\psi:E\rightarrow E_{0}$ with kernel $\sigma(C)$. But having two cyclic $N$-isogenies from $E$ to $E_{0}$ with distinct kernels can happen only if $E$ has complex multiplication (see the proof of [11, Proposition 3.2] for details). We have reached a contradiction. ∎
@@ 3.2. General divisibility conditions
Let $\mathcal{E}$ be rational $\overline{\mathbb{Q}}$-isogeny class of non-CM elliptic curves. Fix a prime number $\ell$ and positive integer $k$. If $E\in\mathcal{E}$, then we can relate the degree of $[E,P]\in X_{1}(\ell^{k})$ to the degree of $[E_{0},P_{0}]\in X_{1}(\ell^{k})$ for some $E_{0}\in\mathcal{E}$ having $j(E_{0})\in\mathbb{Q}$. This is formalized in the following proposition, which strengthens [8, Lemma 4.6] and proves Proposition 1.5.
@@
Let $E\in\mathcal{E}$, and fix $P\in E$ of order $\ell^{k}$. Define $F\coloneqq\mathbb{Q}(j(E),\mathfrak{h}(P))$. By Lemma 3.1, there is a Weierstrass model of $E/F$ where $P\in E(F)$ and such that there exists an $F$-rational cyclic isogeny $\varphi:E\rightarrow E^{\prime}$ with $j(E^{\prime})\in\mathbb{Q}$. By [8, Lemma 4.6], we have $[F:\mathbb{Q}]$ is divisible by
@@
By [8, Corollary 4.3], the curve $E^{\prime}$ has a point of order $\ell$ over an extension $F^{\prime}/F$ of degree dividing $\ell$. In particular, there exists a closed point $x=[E^{\prime},P^{\prime}]\in X_{1}(\ell)$ such that $\mathbb{Q}(x)\subseteq F^{\prime}$. Hence
@@
Since $j(E^{\prime})=j(E_{0})$, there exists a point $P_{0}\in E_{0}$ such that the closed point $x=[E_{0},P_{0}]$. If $\deg(x)$ is prime to $\ell$, then $\deg(x)\mid[F:\mathbb{Q}]$, and the conclusion follows. So suppose $\deg(x)=\ell\cdot x_{0}$. Note $\ell\nmid x_{0}$, since $\deg(X_{1}(\ell)\rightarrow X_{1}(1))<\ell^{2}$, so $x_{0}\mid[F:\mathbb{Q}]$. If there exists $x_{1}\in X_{1}(\ell)$ associated to $E_{0}$ with $\deg(x_{1})\mid x_{0}$, then the conclusion follows with $x_{1}$ in place of $x$. So suppose not. By checking the possible images of the mod $\ell$ Galois representation associated to $E_{0}$ as in [23, Theorem 5.6, Tables 1 & 2], we see that we must be in one of the following cases:
[…]
In each case, there exists a $\mathbb{Q}$-rational cyclic subgroup $C$ of $E_{0}$ such that $E_{0}/C$ has mod $\ell$ image outside this list; see [39, Theorem 3.32, Tables 3 & 4]. That is, in each case there exists a point on $X_{1}(\ell)$ associated to $E_{0}/C$ of degree dividing $x_{0}$, and the result holds with $E_{0}/C$ in place of $E_{0}$. ∎
@@
[…]

### v2

@@ Acknowledgements
We thank John Cremona, Álvaro Lozano-Robledo, Jeremy Rouse, Parker Schwartz, and Andrew Sutherland for helpful conversations. We thank John Cremona, Michael Eddy, Tyler Genao, and Filip Najman for helpful comments on an earlier version of this paper. We are especially grateful to the anonymous referee, who made many comments and suggestions with greatly improved the exposition. All authors were supported by NSF grant DMS-2137659. The first author was partially supported by an A. J. Sterge Faculty Fellowship and NSF grant DMS-2145270.
@@ 2. Background and Notation
Throughout, $F$ denotes a number field and $\overline{F}$ denotes a fixed algebraic closure of $F$. We write $\operatorname{Gal}_{F}$ for the absolute Galois group $\operatorname{Gal}(\overline{F}/F)$.
@@
For a prime number $\ell$, our notation for subgroups of $\operatorname{GL}_{2}(\mathbb{Z}/\ell\mathbb{Z})$ and $\operatorname{GL}_{2}(\mathbb{Z}_{\ell})$ follows [sutherland] and [RSZ21], respectively. For $\ell$-adic images, this is known as the “RSZB label” and has the form N.i.g.n, where N is the level, i is the index, g is the genus, and n is a positive integer used to distinguish nonconjugate subgroups. We also refer to specific elliptic curves over $\mathbb{Q}$ by their $L$-functions and Modular Forms Database [LMFDB] (LMFDB) label.
@@ 2.2. Galois Representations
which encodes the Galois action on all points in $E(\overline{F})$ of order a power of $\ell$. If $E$ is non-CM, then $\operatorname{im}\rho_{E,\ell^{\infty}}$ is an open subgroup of $\operatorname{GL}_{2}(\mathbb{Z}_{\ell})$ by Serre’s Open Image Theorem [serre72]. Thus there exists a nonnegative integer $d$ such that $\operatorname{im}\rho_{E,\ell^{\infty}}=\pi^{-1}(\operatorname{im}\rho_{E,\ell^{d}})$, where $\pi:\operatorname{GL}_{2}(\mathbb{Z}_{\ell})\rightarrow\operatorname{GL}_{2}(\mathbb{Z}/\ell^{d}\mathbb{Z})$ is the natural reduction map. The smallest such $\ell^{d}$ for which this holds is called the level of the $\ell$-adic Galois representation.
@@
Aside from $\operatorname{GL}_{2}(\mathbb{Z}/\ell\mathbb{Z})$, all groups which are known to occur as the image of the mod $\ell$ Galois representation associated to a non-CM elliptic curve over $\mathbb{Q}$ appear (up to conjugacy) in Tables 3 and 4 of [sutherland]. This list is complete for $\ell\leq 13$ by work of Zywina [ZywinaImages] and Balakrishnan, Dogra, Müller, Tuitman, and Vonk [Balakrishnan], and it has been conjectured to be complete for all $\ell$ by both Sutherland [sutherland, Conjecture 1.1] and Zywina [ZywinaImages, Conjecture 1.12]. Unconditionally, we have the following result for $\ell\geq 17$.
Theorem 2.1 (Mazur [mazur78], Serre [Serre81], Bilu, Parent, and Rebolledo [BPR13]).
@@
Suppose $E/\mathbb{Q}$ is a non-CM elliptic curve and $\ell\geq 17$ is prime. If $\operatorname{im}\rho_{E,\ell}$ is not equal to $\operatorname{GL}_{2}(\mathbb{Z}/\ell\mathbb{Z})$ and not conjugate to a group in Table 3 or 4 of [sutherland], then $\operatorname{im}\rho_{E,\ell}$ is contained in $C_{ns}^{+}(\ell)$, the normalizer of a non-split Cartan subgroup of $\operatorname{GL}_{2}(\mathbb{Z}/\ell\mathbb{Z})$.
@@
Refinements of Theorem 2.1 appear in [ZywinaImages, Proposition 1.13], which is also proven in [LeFournLemos, Appendix B]. Recently, Furio and Lombardo [FurioLombardo23] have shown that proper subgroups of the normalizer of a non-split Cartan subgroup do not occur for primes larger than 37. Taken together with prior work, [FurioLombardo23, Theorem 1.6] implies the following result.
Theorem 2.2 (Furio, Lombardo [FurioLombardo23]).
@@
Suppose $E/\mathbb{Q}$ is a non-CM elliptic curve and $\ell\geq 17$ is prime. If $\operatorname{im}\rho_{E,\ell}$ is not equal to $\operatorname{GL}_{2}(\mathbb{Z}/\ell\mathbb{Z})$ and not conjugate to a group in Table 4 of [sutherland], then $\operatorname{im}\rho_{E,\ell}$ is conjugate to $C_{ns}^{+}(\ell)$.
@@
Many partial classification results exist for the image of the $\ell$-adic Galois representation of an elliptic curve $E/\mathbb{Q}$. First suppose $E$ is non-CM. The groups which occur as $\operatorname{im}\rho_{E,2^{\infty}}$ are known due to work of Rouse and Zureick-Brown [RouseDZB]. For odd primes $\ell$, Sutherland and Zywina [SutherlandZywina] have identified the images that occur infinitely often, and work of Rouse, Sutherland, and Zureick-Brown [RSZ21] provides additional classification results for $3\leq\ell\leq 11$ which are complete up to computing rational points on 6 remaining modular curves. If $E$ has complex multiplication, see work of Lozano-Robledo [LRcm].
@@ 2.3. Elliptic Curves with an Isogeny
If the image of $\chi$ lands in $\{\pm 1\}$, then there is a twist of $E$ for which the point corresponding to $P$ becomes $F$-rational. In general, we have the following proposition.
@@
Let $N\geq 3$ be an integer, and let $E/F$ be an elliptic curve with an $F$-rational cyclic isogeny of degree $N$. There is an abelian extension $L/F$ with $L:F]\mid\frac{\varphi(N)}{2}$ and a quadratic twist $E^{\prime}$ of $E/L$ such that $E^{\prime}(L)$ has a point of order $N$.
@@
This is a consequence of [BCS, Theorem 5.5]. ∎
@@ 2.4. Modular Curves
In this paper, we are interested in characterizing degrees of points on the modular curve $X_{1}(N)$, where $N$ is a positive integer. Recall $X_{1}(N)$ is an algebraic curve over $\mathbb{Q}$ whose non-cuspidal points correspond to isomorphism classes of elliptic curves with a distinguished point of order $N$. If $E$ is an elliptic curve defined over a number field $F$ with $P\in E(F)$ of order $N$, then $(E,P)$ gives an $F$-valued point on $X_{1}(N)$ via this moduli interpretation. By definition, this is a morphism of $\mathbb{Q}$-schemes $f:\text{Spec}\,F\rightarrow X_{1}(N)$, and the image of $f$ is the associated closed point, denoted $[E,P]$. See [modular, Section 7.7], [DiamondIm], [shimura, $\S 6.7$], [silverman, Appendix C, $\S 13$], or [DR73] for more details. If $x\in X_{1}(N)$ is closed point, we define the degree of $x$ to be the degree of the residue field $\mathbb{Q}(x)$. For a non-cuspidal point $x$, we can construct $\mathbb{Q}(x)$ explicitly via the following result.
@@
See, for example, [OddDegQCurve, Lemma 2.5], and [DR73, p. 274, Proposition VI.3.2]. ∎
@@
If $E/\mathbb{Q}(j(E))$ corresponds to an equation of the form $y^{2}=x^{3}+Ax+B$ and $P=(x,y)\in E$, then we may take
@@
Note $AB\neq 0$ if $E$ is non-CM. Thus by Lemma 2.4 we can compute the degree of a closed point on $X_{1}(N)$ associated to a non-CM elliptic curve by factoring division polynomials. See [shimura, p. 107] for details, including a formulation of the Weber function which is more clearly model-independent.
@@
The moduli interpretation ensures the map is defined over $\mathbb{Q}$, and the degree calculation follows from [modular, p.66]. ∎
@@ 2.5. Complex Multiplication
where $\Delta_{K}$ is the discriminant of $K$. We have $\mathcal{O}^{\times}=2$ unless $\Delta=-3$ or 4-4 in which case $\mathcal{O}=6$ or $4$, respectively. If $E$ has CM by the order $\mathcal{O}$ in $K$, then $\mathbb{Q}(j(E)):\mathbb{Q}]=h(\mathcal{O})$ by [cox, Theorem 11.1], the class number of $\mathcal{O}$. If $\mathfrak{f}=1$ then $h(\mathcal{O})=h_{K}$, the class number of $K$. If $\mathfrak{f}>1$, then by [cox, Corollary 7.24]
(1) $h(\mathcal{O})=[\mathbb{Q}(j(E)):\mathbb{Q}]=h_{K}\frac{\mathfrak{f}}{[\mathcal{O}_{K}^{\times}:\mathcal{O}^{\times}]}\prod_{p\mid\mathfrak{f}}\left(1-\left(\frac{\Delta_{K}}{p}\right)\frac{1}{p}\right).$
@@ 3.1. Fields of Definition for Isogenies
Let $\mathcal{E}$ be a rational $\overline{\mathbb{Q}}$-isogeny class of non-CM elliptic curves. By definition, there exists $E_{0}\in\mathcal{E}$ with $j(E_{0})\in\mathbb{Q}$, and for any $E\in\mathcal{E}$ there is a $\overline{\mathbb{Q}}$-isogeny $\varphi:E\rightarrow E_{0}$. Since the degree of closed points on $X_{1}(N)$ can be computed using any Weierstrass model of $E$ by Lemma 2.4, we are free to replace $E$ and $E_{0}$ by quadratic twists in order to achieve a more convenient representation of $\varphi$. The lemma given below essentially follows from the fact that $\mathbb{Q}(j(E),j(E_{0}))=\mathbb{Q}(j(E))$ is contained in the residue field of any closed point on $X_{1}(N)$ associated to $E$, and this is the field of moduli of the isogeny $\varphi$; see [ClarkVolcanoes, $\S 3.3$] or [CremonaNajmanQCurve, Corollary A.5].
@@
Since $\mathcal{E}$ is rational, there exists $E_{0}\in\mathcal{E}$ with $j(E_{0})\in\mathbb{Q}$. By definition there exists an isogeny $\varphi:E\rightarrow E_{0}$ defined over $\overline{\mathbb{Q}}$ which we may assume is cyclic of degree $N$; see Lemma A.1 in [CremonaNajmanQCurve]. Let $C$ denote its kernel. Note that $F=\mathbb{Q}(j(E),\mathfrak{h}(P))$ by Lemma 2.4 and there exists a Weierstrass equation of $E/F$ with $P\in E(F)$. The proof strategy of [ClarkVolcanoes, Proposition 3.2] shows $C$ is $F$-rational, as we will now explain. Suppose $\sigma(C)\neq C$ for some $\sigma\in\operatorname{Gal}_{F}$, and consider the induced isogeny $E^{\sigma}\rightarrow(E/C)^{\sigma}$. Since $j(E/C)=j(E_{0})\in\mathbb{Q}$, we see that $j((E/C)^{\sigma})=j(E_{0})$. Thus composition with an isomorphism to $E_{0}$ yields a cyclic $N$-isogeny $\psi:E\rightarrow E_{0}$ with kernel $\sigma(C)$. But having two cyclic $N$-isogenies from $E$ to $E_{0}$ with distinct kernels can happen only if $E$ has complex multiplication (see the last paragraph of the proof of [ClarkVolcanoes, Proposition 3.2] for details). We have reached a contradiction. ∎
@@ 3.2. General divisibility conditions
Let $\mathcal{E}$ be rational $\overline{\mathbb{Q}}$-isogeny class of non-CM elliptic curves. Fix a prime number $\ell$ and positive integer $k$. If $E\in\mathcal{E}$, then we can relate the degree of $E,P]\in X_{1}(\ell^{k})$ to the degree of $E_{0},P_{0}]\in X_{1}(\ell)$ for some $E_{0}\in\mathcal{E}$ having $j(E_{0})\in\mathbb{Q}$. This is formalized in the following proposition, which strengthens [OddDegQCurve, Lemma 4.6] and proves Proposition 1.5.
@@
Let $E\in\mathcal{E}$, and fix $P\in E$ of order $\ell^{k}$. Define $F\coloneqq\mathbb{Q}(j(E),\mathfrak{h}(P))$. By Lemma 3.1, there is a Weierstrass model of $E/F$ where $P\in E(F)$ and such that there exists an $F$-rational cyclic isogeny $\varphi:E\rightarrow E^{\prime}$ with $j(E^{\prime})\in\mathbb{Q}$. By [OddDegQCurve, Lemma 4.6], we have $F:\mathbb{Q}]$ is divisible by
@@
By [OddDegQCurve, Corollary 4.3 (2)], the curve $E^{\prime}$ has a point of order $\ell$ over an extension $F^{\prime}/F$ of degree dividing $\ell$. In particular, there exists a closed point $x=[E^{\prime},P^{\prime}]\in X_{1}(\ell)$ such that $\mathbb{Q}(x)\subseteq F^{\prime}$. Hence
@@
Since $j(E^{\prime})=j(E_{0})$, there exists a point $P_{0}\in E_{0}$ such that the closed point $x=[E_{0},P_{0}]$. If $\deg(x)$ is prime to $\ell$, then $\deg(x)\mid[F:\mathbb{Q}]$, and the conclusion follows since $F$ is the residue field of the closed point associated to $[E,P]$ by Lemma 2.4. So suppose $\deg(x)=\ell\cdot n_{0}$. Note that $\ell\nmid n_{0}$, since $\deg(X_{1}(\ell)\rightarrow X_{1}(1))<\ell^{2}$, so $n_{0}\mid[F:\mathbb{Q}]$. If there exists $x_{1}\in X_{1}(\ell)$ associated to $E_{0}$ with $\deg(x_{1})\mid n_{0}$, then the conclusion follows with $x_{1}$ in place of $x$. So suppose not. By checking the possible images of the mod $\ell$ Galois representation associated to $E_{0}$ as in [GJNajman, Theorem 5.6, Tables 1 & 2], we see that we must be in one of the following cases (see Section 2.1 for an explanation of subgroup labels used):
  • $\ell=5$, $\operatorname{im}\rho_{E_{0},5}=5B.1.2,\,5B.1.3,\text{ or }5B.4.2$,
  • $\ell=7$, $\operatorname{im}\rho_{E_{0},7}=7B.1.3,\,7B.1.4,\text{ or }7B.6.3$,
  • $\ell=13$, $\operatorname{im}\rho_{E_{0},13}=13B.3.2,\,13B.3.7,\,13B.5.2,\text{ or }13B.4.2$,
  • $\ell=17$, $\operatorname{im}\rho_{E_{0},17}=17B.4.6$,
  • $\ell=37$, $\operatorname{im}\rho_{E_{0},37}=37B.8.2$.
In each case, there exists a $\mathbb{Q}$-rational cyclic subgroup $C$ of $E_{0}$ such that $E_{0}/C$ has mod $\ell$ image outside this list; see [sutherland, Theorem 3.32, Tables 3 & 4]. That is, in each case there exists a point on $X_{1}(\ell)$ associated to $E_{0}/C$ of degree dividing $n_{0}$, and the result holds with $E_{0}/C$ in place of $E_{0}$. ∎
@@
  • $\ell=5$, $\operatorname{im}\rho_{E_{0},5}\neq 5B.1.2,\,5B.1.3,\text{ or }5B.4.2$,
  • $\ell=7$, $\operatorname{im}\rho_{E_{0},7}\neq 7B.1.3,\,7B.1.4,\text{ or }7B.6.3$,
  • $\ell=13$, $\operatorname{im}\rho_{E_{0},13}\neq 13B.3.2,\,13B.3.7,\,13B.5.2,\text{ or }13B.4.2$,
  • $\ell=17$, $\operatorname{im}\rho_{E_{0},17}\neq 17B.4.6$,
  • $\ell=37$, $\operatorname{im}\rho_{E_{0},37}\neq 37B.8.2$.